# First-Order Vortex Lattice Melting in Bilayer Ice: A Monte Carlo Method Study

### v1

@@ Abstract
Inspired by the stable bilayer water ice grown in the laboratory (Nature 577, 60, 2020), we propose a model representing water ice as a two-layer six-vertex model. Using the loop update Monte Carlo method, we unveil meaningful findings. While the square lattice six-vertex model exhibits an antiferromagnetic to disordered phase transition known as the Berezinskii-Kosterlitz-Thouless transition, we observe a different scenario for the bilayer six-vertex model, where the transition type transforms into an Ising transition. We discover the emergence of vortices in the disordered phase, and to stabilize them, vortex excitation is induced. This leads to the presence of distinct 1/2 filling and 2/3 filling vortex lattice phases. Importantly, we identify the phase transitions between the vortex lattice phase and the disorder phase, as well as between the 1/2 and 2/3 vortex lattices, as being of first order. Our findings provide valuable insights into the nature of phase transitions occurring in layered water ice and artificial spin ice systems in experimental setups.
@@
Ice is a common substance in nature. There are various types of ice, including the solid form of liquid water [1], spin ice in real materials [2], and artificial spin ice [3, 4]. One common feature of the different forms of ice is the ice rule, i.e., the so called two-in (close) two-out (far away) topological constraint, introduced by Pauling in 1935 [5]. Water ice exhibits 19 stable geometric structures, currently identified through high pressure and low temperature experiments [1]. Spin ice also exists in a lot to real materials with different structures, such as rare-earth pyrochlores $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$ [6].
@@
Researchers have also attempted to grow artificial spin ices  [7], due to their controllability. The microscale systems used to create artificial spin ice typically involve magnetically interacting nanoislands or nanowire links[8], superconducting-qubit arrays [9].The physics studied through ice are very wide, such as residual entropy  [10], frustration [2], monopoles [11] and so on.
@@
Artificial spin ice can construct various configurations, including vortex lattice (VL) phase [12]. The periodic structure composed of vortices is called the VL. As vortex states can help understand superconductors [13], a lot of studies have been done on transitions between vortex lattices and other phases [14, 15, 16, 17]. In the XY model, the unbinding of vortex-antivortex pairs is considered the cause of Berezinskii-Kosterlitz-Thouless (BKT) phase transition [18, 19, 20]. However, the vortex lattice leads to a first-order phase transition in real crystals YBCO and $\text{YBa}_{2}\text{Cu}_{3}\text{O}_{7}$  [14, 15, 16, 17].
@@
Of course, the ice physics can also be explored through experiments with water ice. In 2020, the group at Peking University has confirmed the existence of two-dimensional two-layered water ice [21]. The positions of the upper layer oxygen ions, as well as the connections between the oxygen ions, are exactly the same as those in the lower layer, i.e., AA stacking ice. Such stable structure of water was first predicted in 1997 using molecular dynamics simulation [22]. Actually, there exists another type of ice with an AB stacking structure [23] as shown Fig. 1. There is a relative 180 degree rotation between the two layers, which are connected by hydrogen bonds. Alternatively, the hydrogen bonds of the first layer are shifted onto the faces of the second layer. It has been confirmed to be stable under reasonable temperature and pressure by first-principles calculations [23] but lack of experimental preparation [21].
@@
Figure 1: The construction of the two-layer ice model. The large balls (in red and blue) represent $O^{2-}$s and the small balls denote $H^{+}$s. The different colors for the large ball means AB stacking [23]. The two-in-two-out rule can be demonstrated by the small balls in different colors.
@@
To inspire experimental physicists to achieve AB or other type of bilayer ice in the future, we further convert bilayer ice to bilayer six-vertex (6V) model model [24]. On the basis of the bilayer square lattice 6V model, the AB stacking honeycomb 6V model can be formed by bending one leg of the vertexes to the appropriately. Hydrogen ions and oxygen ions in ice can be close or far apart, similar to spin-up and spin-down. Viewing ice as a spin system helps understand the phase transitions between different types of ice under various conditions like temperature and pressure. It is not clear that whether or not there are something new in the bilayer honeycomb 6V model.
@@
In this paper, we apply a large-scale loop Monte Carlo (MC) simulation to study the proposed 6V model. By properly defining and scanning the type and weight, including the vortex-exciting weight, we explore the phase diagram systematically. The system includes ferromagnetic, antiferromagnetic, 1/2-filling VL, 2/3-filling VL, and disordered phases. The theoretically discovered types of phase transitions, such as first-order phase transitions, also provides insights into understanding previous experiments [14, 15, 16, 17].
@@
The outline of this work is as follows. Sec. II introduces the bilayer 6V model, algorithm, and the measured quantities. Sec. III describes the phase diagram and details without the effects of vortex weight $V=1$. Sec. IV describes the phase diagram with the effects of vortex weight $V\neq 1$. Conclusive comments and outlook are made in Sec. V.
@@
Physically, apart from the bilayer 6V model initially propose, we have made new discoveries as follows:
@@
(I) The transition between the antiferromagnetic phase to the disordered phase is the Ising type for our bilayer 6V model. However, for the 6V model on the square lattice, the transition is the BKT type [25].
@@
(II) Two types of vortex lattice phase is found when vortex excitation is induced. The transition from vortex lattice phases to other phases is of first order, consistent with previous crystal experiments [14, 15] . It helps researchers understand that not all phase transitions involving vortices are BKT phase transitions.
@@
On the algorithmic level, although Ref. [24] has simulated the single-layer planar 6V model using loop algorithm, the model here is the non-planar 6V model and we provide the details of closing loops. In addition, we introduce a Metropolis-type of short-loop update method for the purpose of ergodicity.
@@
Figure 2: Schematic diagram of the 6V model and water molecules. (a1)-(c2) The configurations of six vertices and their weights defined as $a$, $a$, $b$, $b$, $c$ and $c$, respectively. (d1)-(d2) Connecting the two vertices by the two vertical red legs marked by 2 and 4. (e1)-(e2) Rotating the two legs, blue respectively, on the horizontal plane. (f1)-(f2) Mapping to the structure of real water ice molecules. The small ball represents hydrogen ion and the large ball represents oxygen ion. (g1)-(g2) The definition of vortex.
@@ II.1 The 6V model with vortex weight
II.1.1 Hamilton and partition function
Unlike the models such as the Ising [26], XY [18, 19, 20], Potts  [27] models or the coupled spins such Ising-XY model [28], etc., the famous 6V model does not have a explicit Hamiltonian. But each type of vertex has its own weight and can also have an equivalent energy, and the vertex satisfies the two-in-two-out topological constraints. For convenience, to measure physical quantities related to specific heat and other energy-related quantities, a quasi-Hamiltonian is introduced as
@@ II.1.1 Hamilton and partition function
Using the Boltzmann weight factor, the vertex weight is
@@
where $\beta=1/k_{B}T$ is the inverse temperature and set to 1, and the partition function of the system is defined as follows
@@
Figures 2 (a1)-(c2) show the configurations of six vertexes, and each vertex has four legs marked by the red arrows. The directions of the legs satisfy two facing out away from the center and two legs pointing toward the center. For simplicity, the weights of the six vertexes take values as
$w_{i}=[a,a,b,b,c,c]$ (5)
here the flipping symmetry of the leg direction is considered, i.e., the state of (a1) can be obtained by flipping the directions of the four legs of (a2), and therefore the weights of these two vertexes are the same as $a$. Totally, only three possible weight values $a$, $b$ and $c$ are needed for the six different configurations. This tradition is also used in Ref. [25].
It should be note that the vortex is defined as shown in Figs. 2 (g1) and (g2). Numerically, the requirement is that the angle difference between adjacent vectors is $\frac{\pi}{3}$, i.e.,
@@
where $i$ ranges from 1 to 6. This type of definition does not need the saw function in Ref. [28] and is the same the one in Ref. [29].
@@ II.1.2 The formation of two-layer honeycomb lattice structure
Figures 2(d1)-(e2) show how to use the vertexes to construct a two-layer honeycomb lattices. Firstly, 6V models are usually constructed in the two-dimensional plane, and it is rare to see two-layer 6V models. Here, the interlayer coupling is realized by connecting the two vertices, using the two red legs marked 2, 4 in the vertical direction. Then, the second step is rotating the two legs in blue to the horizontal $y$-direction. A small coordinate axes $x$, $y$, $z$ are shown for reference.
It is also essential to illustrate the relationship between the two-layer honeycomb 6V model and the structure of ordered water molecules. In Figs. 2(f1)-(f2), the arrows pointing towards the center of the vertex means, the approaching of the hydrogen (H) ions to the oxygen (O) ion in a real water molecule. Conversely, arrows pointing away from the center indicate that the hydrogen ions stays away from the oxygen ion. Ultimately, the arrangement of vertices can describe the structure of real water ice molecules.
In addition, the configurations of the system also depends the topological constraints of “two-in and two-out” rules [24, 30]. In other words, the arrows in the vertices have two pointing to the center and the other two back to the center. The real configurations of large lattices needs to be simulated by various MC methods [31, 32], which we will discussed in next section.
Figure 3: Schematic diagram of the loop algorithm in a two-layer honeycomb lattice. The start point is selected randomly, and the red arrows illustrate the path of the process. When the head of the loop meet the start vertex, the loop is closed and the configuration is updated. For convenience in defining the coordinates of the vertices, the honeycomb is defined as a topologically equivalent brick wall structure.
@@ II.2.1 Methods
In this paper, we apply the loop algorithm, which has has proven effective in studying various systems both classical systems systems [24], and quantum systems [33, 34]. A similar loop algorithm is the famous worm algorithm [35, 36], which involves a partial loop with two open ends with very efficient dynamical behaviours [37]. The 6V model is very similar to the flow representation of other models [38, 39, 40].
@@
Let us explain how the probability of selecting the exit leg when the loop head is determined. Here, we employ the Metropolis-Hastings strategy. Suppose the weight of the reference vertex is $w_{0}$, and the weights of the new vertices resulting from exiting from the four legs are $w_{1}$, $w_{2}$, $w_{3}$, and $w_{4}$ respectively. We choose a random number between 0 and 1 to determine the interval in which the random number falls. These intervals are defined as $[0,p_{1}]$, $(p_{1},p_{1}+p_{2}]$, $(p_{1}+p_{2},p_{1}+p_{2}+p_{3}]$, and $(p_{1}+p_{2}+p_{3},1]$, where the probabilities are defined as:
@@
If the outgoing and incoming legs are exactly the same, then there is no update, and the corresponding probability $p$ is referred to as the bounce probability. Generally, a higher bounce probability leads to lower efficiency of the cluster algorithm [33, 34].
Therefore, analyzing the possible range of bounce probability under different parameters (see Table 1) here would be helpful to ensure the feasibility of the code.
@@
Loop close is an important step in the loop algorithm. In some cases, the length of the loop is very big and the code consumes very long runtimes. Ref. [41] even cut the loop by the so-called short-loop mehtods.
Here, we resort to the method dealing with the quantum Bose-Hubbard models [42, 43, 44, 45], two ways of closing loops are used. In Fig. 4 (a), the beginning leg labeled by “first”, the ending leg is marked by “last”. In this case, the first and the last legs meet (overlap) at the same leg, and then the loop close. In Fig. 4 (b), the last leg connects to the position of the initial let. This constraint arises from the “2-in and 2-out” condition, allowing the loop closure only when connected to the initial leg. The distinction lies in the fact that in the former case, the vertex labeled “1” undergoes two updates, whereas in the latter, it undergoes only one time of update.
Figure 4: The two ways in which the loops close. (a) First and the last leg overlap at the same leg, and (b) the last leg connecting the first leg but the two leg belong to a pair of neighborhood Vertexes.
@@ II.2.2 The measured quantities
where (±$(\pm 1)$ mean the factor of phase of ferro-magnetization (FM) or antiferro-magnetization (AFM) respectively. The symbols $x$ refer to the horizontal directions of the lattice, and $i$ and $j$ means the coordinates of the vertices in the $x$ and $y$ directions.
(Ib) The absolute values of magnetizations are
@@
(Ic) The striped ferro-magnetization defined as
$m_{s}=\sum_{i=1}^{L}\left|\sum_{j=1}^{L}S_{i,j}\right|/N,$ (10)
which is further to defined the striped specific heat $C_{ms}$.
(Id) The Binder ratio is also defined as
@@
$Q_{s}$ $={\langle m_{s}^{2}\rangle}^{2}/{\langle{(m_{s}}^{4}\rangle}$ (12)
@@
where $f$ is the number of total faces of one layer.
(III) Specific heats $C_{V}$, and $C_{vx}$ are expressed as
@@
where $e$ is the average energy persite. $C_{V}$, and $C_{vx}$ are specific heats related with energy, and vortex.
III phase diagram $V=1$ ($v=0$) and details
III.1 global phase diagram with $V=1$

### v2

@@ Abstract
Inspired by the stable bilayer water ice grown in the laboratory (Nature 577, 60, 2020), we propose a model representing water ice as a two-layer six-vertex model. Using the loop update Monte Carlo method, we unveil meaningful findings. While the square lattice six-vertex model exhibits an antiferromagnetic to disordered phase transition known as the Berezinskii-Kosterlitz-Thouless transition, we observe a different scenario for the bilayer six-vertex model, where the transition type transforms into an Ising transition. We discover the emergence of vortices in the disordered phase, and to stabilize them, vortex excitation is induced. This leads to the presence of distinct 1/2 filling and 2/3 filling vortex lattice phases. More importantly, we identify the phase transitions between the vortex lattice phase and the disordered phase, as well as between the 1/2 and 2/3 vortex lattices, as being of first order. We also propose an experimental scheme for realizing a two-layer six-vertex model based on the artificial ice of particles in a double well trap array. Our findings provide valuable insight into the nature of phase transitions occurring in layered water ice and artificial spin ice systems in experimental setups.
@@
Ice is a common substance in nature. There are various types of ice, including the solid form of liquid water [1], spin ice in real materials [2], artificial spin ice [3, 4], and particle ice [5]. One common feature of the different forms of ice is the ice rule, i.e., the so called two-in (close) two-out (far away) topological constraint, introduced by Pauling in 1935 [6]. Water ice exhibits 19 stable geometric structures, currently identified through high pressure and low temperature experiments [1]. Spin ice also exists many real materials with different structures, such as rare-earth pyrochlores $\text{Ho}_{2}\text{Ti}_{2}\text{O}_{7}$ [7].
@@
Researchers have also attempted to grow artificial spin ices [8], due to their controllability. The microscale systems used to create artificial spin ice typically involve magnetically interacting nanoislands or nanowire links[9], superconducting-qubit arrays [10].The physics studied through ice is very broad; it covers residual entropy  [11], frustration [2], monopoles [12] and so forth.
@@
Artificial spin ice can construct various configurations, including the vortex lattice (VL) phase [13]. Its periodic structure composed of vortices is called the VL. As vortex states can help understand superconductors [14], many studies have been done on transitions between vortex lattices and other phases [15, 16, 17, 18]. In the XY model, the unbinding of the vortex-antivortex pairs is considered the cause of the Berezinskii-Kosterlitz-Thouless (BKT) phase transition [19, 20, 21]. However, the vortex lattice leads to a first order phase transition in the real crystals YBCO and $\text{YBa}_{2}\text{Cu}_{3}\text{O}_{7}$  [15, 16, 17, 18].
@@
Of course, the ice physics can also be explored through experiments with water ice. In 2020, the group at Peking University confirmed the existence of two-dimensional two-layered water ice [22]. The positions of the upper layer oxygen ions, as well as the connections between the oxygen ions, are exactly the same as those in the lower layer, i.e., AA stacking ice. Such a stable structure of water was first predicted in 1997 using a molecular dynamics simulation [23]. Actually, there exists another type of ice with an AB stacking structure [24] as shown Fig. 1. There is a relative 180 degree rotation between the two layers, which are connected by hydrogen bonds. Alternatively, the hydrogen bonds of the first layer are shifted onto the faces of the second layer. It has been confirmed to be stable under reasonable temperatures and pressures by ab-initio calculations [24] though lacking in experimental preparation [22].
@@
Figure 1: The construction of the two-layer ice model. The large balls (in red and blue) represent $O^{2-}$s and the small balls denote $H^{+}$s. The different colors for the large ball means AB stacking [24]. The “two-in-two-out rule can be demonstrated by the small balls in different colors.
@@
To inspire experimental physicists to achieve AB or other types of bilayer ice in the future, we further convert the bilayer ice to the bilayer six-vertex (6V) model [25]. On the basis of the bilayer square lattice 6V model, the AB stacking honeycomb 6V model can be formed by bending one leg of the vertices to the appropriate degree. Hydrogen ions and oxygen ions in ice can be close to each other or far apart, similar to the situation of spin-up and spin-down. Viewing ice as a spin system helps understand the phase transitions between different types of ice under various conditions like temperature and pressure. It is not clear whether or not there is something new in the bilayer honeycomb 6V model.
@@
In this paper, we apply a large-scale loop Monte Carlo (MC) simulation to study the proposed 6V model. By properly defining and scanning the type and weight, including the vortex exciting weight, we explore the phase diagrams systematically. The system includes ferromagnetic, antiferromagnetic, 1/2-filling VL, 2/3-filling VL, and disordered phases. The theoretically discovered types of phase transitions, such as first order phase transitions, also provide insight into understanding previous experiments [15, 16, 17, 18].
@@
The outline of this work is as follows. Sec. II introduces the bilayer 6V model, algorithm, and the measured quantities. Sec. III describes the phase diagram and details without the effects of vortex weight $V=1$. Sec. IV describes the phase diagram with the effects of vortex weight $V\neq 1$. Sec. V discusses the experimental realization. Conclusive comments and outlook are made in Sec. VI.
@@
Physically, apart from the bilayer 6V model initially proposed, we have made new discoveries as follows:
@@
(I) The transition between the antiferromagnetic phase to the disordered phase is of Ising type for our bilayer 6V model. However, for the 6V model on the square lattice, the transition is of BKT type [26].
@@
(II) Two types of vortex lattice phase are found when vortex excitation is induced. The transition from vortex lattice phases to other phases is of first order, consistent with previous crystal experiments [15, 16] . This helps researchers understand that not all phase transitions involving vortices are BKT phase transitions.
@@
On the algorithmic level, although Ref. [25] has simulated the single-layer planar 6V model using the loop algorithm, our model is the non-planar 6V model and we provide the details of closing loops. In addition, we introduce a Metropolis type short-loop update method for the purpose of ergodicity.
Experimentally, bilayer water ice, where the ordering is due to the proton positions [22], is probably better captured by particle-based ice [5]. Here, we propose to groove or imprint both bilayer lattices onto a single-layer substrate, and each edge of the lattices are realized by a double well trap with only one particle.
@@
Figure 2: Schematic diagram of the 6V model and water molecules. (a1)-(c2) The configurations of six vertices and their weights defined as $a$, $a$, $b$, $b$, $c$ and $c$, respectively. (d1)-(d2) Connecting the two vertices by the two vertical red legs marked by 2 and 4. (e1)-(e2) Rotating the two legs, marked in blue, on the horizontal plane. (f1)-(f2) Mapping to the structure of real water ice molecules. The small ball represents the hydrogen ion and the large ball represents the oxygen ion. (g1)-(g2) The definition of the vortex.
@@ II.1 The 6V model with vortex weight
II.1.1 Hamiltonian and partition function
Unlike the models such as the Ising [27], XY [19, 20, 21], Potts  [28] models or the coupled spins such the Ising-XY model [29], etc., the famous 6V model does not have a explicit Hamiltonian. However, each type of vertex has its own weight and can also have an equivalent energy, and the vertex satisfies the “two-in-two-out” topological constraints. For convenience, to measure physical quantities related to specific heat and other energy-related quantities, a quasi-Hamiltonian is introduced as:
@@ II.1.1 Hamiltonian and partition function
Using the Boltzmann weight factor, the vertex weight is:
@@
where $\beta=1/k_{B}T$ is the inverse temperature and set to 1, and the partition function of the system is defined as follows:
@@
Figures 2 (a1)-(c2) show the configurations of six vertices, and each vertex has four legs marked by the red arrows. The directions of the legs satisfy two of them facing out away from the center and two legs pointing toward the center. For simplicity, the weights of the six vertices take values as:
$w_{i}=[a,a,b,b,c,c]~{}.$ (5)
Here, the flipping symmetry of the leg direction is considered, i.e., the state of (a1) can be obtained by flipping the directions of the four legs of (a2), and therefore the weights of these two vertices are the same as $a$. In total, only three possible weight values $a$, $b$ and $c$ are needed for the six different configurations. This tradition is also used in Ref. [26].
It should be noted that the vortex is defined as shown in Figs. 2 (g1) and (g2). Numerically, the requirement is that the angle difference between adjacent vectors is $\frac{\pi}{3}$, i.e.,
@@
where $i$ ranges from 1 to 6. This type of definition does not need the saw function in Ref. [29] and is the same as the one in Ref. [30].
@@ II.1.2 The formation of two-layer honeycomb lattice structure
Figures 2(d1)-(e2) show how to use the vertices to construct two-layer honeycomb lattices. Firstly, 6V models are usually constructed in the two-dimensional plane, and it is rare to see two-layer 6V models. Here, the interlayer coupling is realized by connecting the two vertices, using the two red legs marked 2, 4 in the vertical direction. Then, the second step rotates the two legs (in blue) to the horizontal $y$-direction. A small coordinate axes system denoted $x$, $y$, $z$ is shown for reference.
It is also essential to illustrate the relationship between the two-layer honeycomb 6V model and the structure of ordered water molecules. In Figs. 2(f1)-(f2), the arrows pointing towards the center of the vertex corresponds to the approaching of the hydrogen (H) ions towards the oxygen (O) ion in a real water molecule. Conversely, arrows pointing away from the center indicate that the hydrogen ions stay away from the oxygen ion. Ultimately, the arrangement of vertices can describe the structure of real water ice molecules.
In addition, the configurations of the system also depend on the topological constraints of “two-in-two-out” rules [25, 31]. In other words, the arrows in the vertices have two pointing of them to the center and the other two back to the center. The real configurations of large lattices need to be simulated by various MC methods [32, 33], which will be discussed in the next section.
Figure 3: Schematic diagram of the loop algorithm in a two-layer honeycomb lattice. The starting point is selected randomly, and the red arrows illustrate the path of the process. When the head of the loop meets the start vertex, the loop is closed and the configuration is updated. For convenience in defining the coordinates of the vertices, the honeycomb is defined as a topologically equivalent brick wall structure.
@@ II.2.1 Methods
In this paper, we apply the loop algorithm, which has proven effective in studying various systems both classical systems systems [25], and quantum systems [34, 35]. A similar loop algorithm is the famous worm algorithm [36, 37], which involves a partial loop with two open ends with very efficient dynamical behaviors [38]. The 6V model is very similar to the flow representation of other models [39, 40, 41].
@@
Let’s explain how to choose the exit leg when the ring head has been determined. Here, we employ the Metropolis-Hastings strategy. Suppose the weight of the reference vertex is $w_{0}$, and the weights of the new vertices resulting from exiting from the four legs are $w_{1}$, $w_{2}$, $w_{3}$, and $w_{4}$ respectively. We choose a random number between 0 and 1 to determine the interval in which the random number falls. These intervals are defined as $[0,p_{1}]$, $(p_{1},p_{1}+p_{2}]$, $(p_{1}+p_{2},p_{1}+p_{2}+p_{3}]$, and $(p_{1}+p_{2}+p_{3},1]$, where the probabilities are defined as:
@@
If the outgoing and incoming legs are exactly the same, then there is no update, and the corresponding probability $p$ is referred to, as the bounce probability. Generally, a higher bounce probability leads to lower efficiency of the cluster algorithm [34, 35].
Therefore, analyzing the possible range of bounce probabilities under different parameters (see Table 1) would be helpful to ensure the feasibility of the code.
@@
Loop close is an important step in the loop algorithm. In some cases, the length of the loop is very big and the code consumes very long run times. The authors of Ref. [42] even cut the loop by the so-called short-loop methods.
Here, we resort to the method dealing with the quantum Bose-Hubbard models [43, 44, 45, 46]. Two ways of closing loops are used. In Fig. 4 (a), the beginning leg is labeled by “first” and the ending leg is marked by “last”. In this case, the first and the last legs meet (overlap) at the same leg, and then the loop closes. In Fig. 4 (b), the last leg connects to the position of the initial leg. This constraint arises from the “2-in and 2-out” condition, allowing the loop closure only when connected to the initial leg. The distinction lies in the fact that in the former case, the vertex labeled “1” undergoes two updates, whereas in the latter, it undergoes only a one time update.
Figure 4: The two ways in which the loops close. (a) The first and last legs overlap at the same leg, and (b) the last leg connecting to the first leg but the two legs belong to a pair of neighborhood vertices.
@@ II.2.2 The measured quantities
where (±$(\pm 1)$ denotes the phase factor of ferromagnetization (FM) or antiferromagnetization (AFM) respectively. The symbol $x$ refers to the horizontal directions of the lattice, and $i$ and $j$ mean the coordinates of the vertices in the $x$ and $y$ directions.
(Ib) The absolute values of magnetizations are:
@@
(Ic) The striped ferromagnetization is defined as:
$m_{s}=\sum_{i=1}^{L}\left|\sum_{j=1}^{L}S_{i,j}^{x}\right|/N,$ (10)
which is used later to define the striped specific heat $C_{ms}$.
(Id) The Binder ratios are also defined as:
@@
$Q_{s}$ $={\langle m_{s}^{2}\rangle}^{2}/{\langle{(m_{s}}^{4}\rangle},$ (12)
corresponding to the antiferro magnetization $M^{-}_{x}$ and the striped ferromagnetization $m_{s}$, respectively.
@@
where $f$ is the total number of faces of one layer.
(III) Specific heats $C_{V}$, and $C_{vx}$ are expressed as:
@@
where $e$ is the average energy per site. $C_{V}$ and $C_{vx}$ are specific heats related to energy and vortex.
III phase diagram $\bf V=1$ ($\bf v=0$) and details
III.1 Global phase diagram with $\bf V=1$